### v1

# Dirac’s variational approach to semiclassical Kramers problem in Smoluchowski limit

Abstract
Kramers escape from a metastable state in the presence of both thermal and quantum fluctuations under strong damping is treated as a thermally activated process in a quantum modified semiclassical potential. Dirac’s time-dependent variational method together with the Jackiw-Kerman function is employed to derive the semiclassical potential.
I Introduction
The problem of escape from metastable states is ubiquitous in physics, chemistry, and biology. A particle can escape from a metastable state owing to thermal fluctuations, or quantum tunneling, or both effects.
The work of Kramers K represents a classic analysis of such problems as a thermally activated processes. He treated the escape process as a Brownian motion of the particle in the metastable state driven by thermal noise. The dynamics was given in terms of some probability density function that satisfies an equation of the Fokker-Planck type, called the Kramers equation G ; R .
In recent years there has been much interest in the study of quantum Kramers problem, which considers both the thermal and quantum fluctuations. This is mainly motivated by the work of Caldeira and Leggett M1 . They studied quantum tunneling in the presence of dissipation at zero temperature. Their motivation was to understand quantum mechanics at the macroscopic level as stimulated by the problem of Schrödinger cat. The method of path-integral was found useful in their work, and has since become the main tool in subsequent works by others. Extension to finite temperatures were given in M2 ; M3 ; M4 . Candidates for such macroscopic quantum phenomena include, for example, Josephson junction and SQUID circuit, single-domain magnet, and liquid Helium-3 and -4 mixture M5 . For reviews of the earlier works, see e.g., QK1 ; QK2 . Soon after, different approaches have been proposed to study the problem, e.g., by Wigner phase space W1 ; W2 , and by quantum Smoluchowski equations A1 ; A2 ; A3 ; S1 ; S2 .
For quantum Kramers problem, one is basically interested in the quantum correction to the thermally activated escape rate over a potential barrier. In this case, a semiclassical approach is usually adopted. Most of the works appear so far make use of the complicated machinery of path integral. This method, while elegant, unfortunately involves continuation into imaginary time in its formalism. As such it is not easy to have an intuitive understanding of the physical process involved. Furthermore, unlike the classical theory of Kramers, which considers the evolution of probability distribution function of the Brownian motion of a particle under a force field, the path integral method deals mainly with the evaluation of quantum partition function of the system interacting with a bath of harmonic oscillators having a certain distribution of frequencies. Different distributions represent different damping characteristics. Thus the result depends very much on the frequency spectrum of the oscillators assumed.
In our view, a simple and more intuitive semiclassical approach has been the time-dependent variational method initiated by Dirac D1 ; D2 . This method gives the effective semiclassical Hamiltonian very directly. Previously we have employed this method to study the phases of the ground states of the quantum Frankel-Kontorova model H1 , and the quantum chaotic behavior of the Hénon-Heiles model H2 . In these work we found it very useful to take the Jackiw-Kerman function JK ; C as the trial wavefuncrtion for variational method.
Here we would like to apply this approach to study the semiclassical Kramers escape problem. The aim is to incorporate quantum effect in the semiclassical way, i.e., to see how quantum effect modifies the potential that trap the metastable state. This amounts to obtaining an effective potential with quantum correction. The escape rate from this new potential is then calculated following Kramers’ original method.
II Classical Kramers escape rate
A particle of mass $M$ moving in a potential $V(x)$ together with a damping force and a stochastic force is governed by the stochastic equation
$M\ddot{x}(t)+M\gamma\dot{x}(t)+V^{\prime}(x)=M\gamma\,\xi(t).$ (1)
Here the “dot” and the “prime” represent derivatives with respect to time and space, respectively. $-M\gamma\dot{x}$ is the damping force with damping rate per unit mass $\gamma$. The stochastic force $\xi(t)$ depends only on the time $t$ and obeys the stationary Gaussian statistics
$\langle\xi(t)\rangle=0,~{}~{}\langle\xi(t)\xi(0)\rangle=2D\delta(t),~{}~{}~{}D\equiv\frac{k_{B}T}{M\gamma},$ (2)
where $k_{B}$ is the Boltzmann constant and $T$ the temperature.
For strong damping, one can ignore the first term in (1) and obtain the Langevin equation
$\dot{x}(t)+\frac{V^{\prime}(x)}{M\gamma}=\xi(t).$ (3)
By Itô’s prescription, the Fokker-Planck equation corresponding to (3) is the Smoluchowski equation
$\frac{\partial}{\partial t}P(x,t)=\frac{\partial}{\partial x}\left(\frac{V^{\prime}(x)}{M\gamma}\,P(x,t)\right)+D\frac{\partial^{2}}{\partial x^{2}}\,P(x,t).$ (4)
Here $P(x,t)$ is the probability density function describing the motion of the particle, and $V^{\prime}(x)/M\gamma$ and $D$ are the drift and the diffusion coefficient, respectively.
From now on we shall consider the escape problem of a metastable state from a potential $V(x)$ having a minimum $V(x_{0})=0$ at the origin $x_{0}=0$ and a maximum with barrier height $V(x_{b})>0$ at $x_{b}>0$, together with the boundary conditions $V(x)\to\infty,V_{\infty}<0$ as $x\to-\infty,\infty$, respectively. The particle, once escape, will move towards $x\to+\infty$. One such potential is the cubic potential, $V(x)=Ax^{2}-Bx^{3}(A,B>0)$. We consider the situation where $V(x_{b})\gg k_{B}T\gg\hbar\omega_{0}$, here $\omega_{0}\equiv\sqrt{V^{\prime\prime}(x_{0})/M}$ is the frequency of small oscillation around the minimum $V(x_{0})$.
By taking quadratic approximation at the minimal and maximal points, Kramers obtains the classical escape rate K ; G ; R
$r_{c}=\frac{\omega_{0}\omega_{b}}{2\pi\gamma}\,e^{-\beta[V(x_{b})-V(x_{0})]}$ (5)
for the over-damped case $\gamma\gg\omega_{b}$, where $\omega_{b}\equiv\sqrt{|V^{\prime\prime}(x_{b})|/M}$ is the frequency around $x_{b}$ of the inverted potential.
We would like to see how quantum fluctuations would modify the above escape rate. To account for quantum fluctuations, we consider a semiclassical model of the escape problem. We modify the potential $V(x)$ to an effective semiclassical potential ${\hat{V}}(x)$ that incorporates quantum effects. The motion of the escaping particle is then considered to be caused by the extra stochastic noise in this new potential. To derive the effective semiclassical potential, we adopt Dirac’s time-dependent variational principle D1 together with the Jackiw-Kerman (JK) function JK . The essence of this approach we shall briefly review in the next section.
III Dirac’s time-dependent variational approach
Quantum Hamiltonian $H$ of a particle of mass $M$ in the potential $V(Q)$ is $H=\frac{P^{2}}{2M}+V(Q)$, where $P,Q$ are the momentum and position operators.
[…]
$\langle Q|\Psi,t\rangle=\frac{1}{(2\pi\hbar G)^{1/4}}$ (6)
×$\times\exp\Biggl{\{}-\frac{1}{2\hbar}\left(Q-x\right)^{2}\Bigl{[}\frac{1}{2}G^{-1}-2i\Pi\Bigr{]}+\frac{i}{\hbar}p\left(Q-x\right)\Biggr{\}}.$
The real quantities $x(t)$, $p(t)$, $G(t)$ and $\Pi(t)$ are variational parameters the variations of which at $t=\pm\infty$ are assumed to vanish. We prefer to use the JK form since the physical meanings of the variational parameters contained in the JK wavefunction are most transparent, as will be evident from the discussion below.
The effective action $\Gamma$ for the Hamiltonian $H$ can be worked out to be C ; H1 ; H2
[…]
where $H_{sc}=\langle\Psi|H|\Psi\rangle$ is the semiclassical Hamiltonian evaluated to be, up to $\hbar$-terms,
$~{}~{}H_{sc}=\frac{1}{2M}p^{2}+V_{sc}(x),$ (8)
$V_{sc}(x)=V(x)+\hbar\left[\frac{1}{8MG}++\frac{2}{M}G\Pi^{2}+\frac{1}{2}GV^{\prime\prime}\right].$
One sees from the form of the effective action $\Gamma$ that $\Pi$ is the canonical conjugate of $\hbar G$. The second line of (8) gives the quantum contribution to the classical Hamiltonian in this semiclassical model.
It is not hard to work out the following expectation values:
$\langle\Psi|Q|\Psi\rangle=x,\quad\quad\langle\Psi|P|\Psi\rangle=p,$ (9)
$\langle\Psi|(Q-x)^{2}|\Psi\rangle=\hbar G,$ (10)
$\langle\Psi|(P-p)^{2}|\Psi\rangle=4\hbar\,G\Pi^{2}+\frac{\hbar}{4G}.$ (11)
It is clear that $x$ and $p$ are the expectation values of the operators $Q$ and $P$, respectively. Also, $\hbar G$ is the mean fluctuation of the position and that $G>0$. $\Pi$ is related to the mean fluctuation of $P$. The uncertainty relation is
$\Delta Q\,\Delta P=\frac{\hbar}{2}\sqrt{1+(4G\Pi)^{2}}.$ (12)
Varying $\Gamma$ with respect to $x,~{}p,~{}G$ and $\Pi$ gives the Hamilton equations of motion
$\dot{x}=\frac{p}{M},\quad\quad{\dot{G}}=\frac{4}{M}G\Pi,$
$\dot{p}=-V^{\prime}-\frac{\hbar}{2}GV^{\prime\prime\prime},$ (13)
$\dot{\Pi}=\frac{1}{8M}G^{-2}-\frac{2}{M}\Pi^{2}-\frac{1}{2}V^{\prime\prime}.$
IV Effective potential and semi-classical Kramers escape rate
To derive the effective potential appropriate for our model of the semiclassical Kramers problem, we make the following assumptions: 1) the quantum fluctuations should be as small as possible, so that the system is as close to the classical limit as possible; 2) in the strong damping limit, it is reasonable to assume the particle stays mainly in the bottom of the well before crossing the barrier.
The requirement in assumption (1) can be attained by requiring that the uncertainly relation between $Q$ and $P$ be minimal. This requires $\Pi=0$ from (12). Note that we must have $G>0$ in order that $V_{sc}(x)$ in (8) be a regular function.
Assumption (2) implies the particle is nearly in a quasi-stationary state at $x_{0}=0$. So we could assume that the state variables vary very slowly. Here we consider the situation in which the quantum fluctuations are nearly time-independent, i.e., $\dot{G}=\dot{\Pi}=0$. That $\dot{G}=0$ is already implied by $\Pi=0$ from the second equation in (13). Likewise we demand $\Pi=0$ is constant in time, i.e. $\dot{\Pi}=0$. The forth equation in (13) then fixes the value of $G$ to be
$G=\frac{1}{2\sqrt{MV^{\prime\prime}(x_{0})}}.$ (14)
Putting these values back into $H_{sc}$ then gives us the effective potential
${\hat{V}}(x)=V(x)+\hbar\left[\frac{1}{8MG}++\frac{1}{2}GV^{\prime\prime}\right],$ (15)
where $G$ is now given by (14). Equation of motion of $p$ in (13) is the same as the Newton second law using ${\hat{V}}(x)$.
We now treat the Kramers problem in the presence of both the thermal and the quantum fluctuations as simply the Kramers escape activated by thermal noise in the semiclassical potential ${\hat{V}}(x)$. The derivation of the escape rate is then carried out in exactly the same way as in the classical case. It is noted that, for strong damping and small quantum fluctuations assumed here, the original potential $V(x)$ is deformed only slightly to $\hat{V}(x)$, and that the quadratic approximation at the minimum and maximum of $\hat{V}(x)$ is still appropriate. Thus the quantum corrected escape rate is given by the expression (5), with all relevant quantities related to $V(x)$ replaced by those related to $\hat{V}(x)$,
$r_{sc}=\frac{\hat{\omega}_{0}\hat{\omega}_{b}}{2\pi\gamma}\,e^{-\beta[\hat{V}(\hat{x}_{b})-\hat{V}(\hat{x}_{0})]}.$ (16)
Here $\hat{x}_{0}$ and $\hat{x}_{b}$ are the positions of the minimum and maximum of $\hat{V}(x)$, and $\hat{\omega}_{0}\equiv\sqrt{\hat{V}^{\prime\prime}(\hat{x}_{0})/M}$ and $\hat{\omega}_{b}\equiv\sqrt{|\hat{V}^{\prime\prime}(\hat{x}_{b})|/M}$.
Again, under the approximation adopted here, we can set $\hat{x}_{0,b}\approx x_{0,b}$ and $\hat{\omega}_{0,b}\approx\omega_{0,b}$. With these values, we finally have
$r_{sc}\approx r_{c}e^{\frac{1}{2}\beta\hbar G\left[V^{\prime\prime}(x_{0})+|V^{\prime\prime}(x_{b})|\right]},$ (17)
where $r_{c}$ is given by (5). The second exponential factor represents the enhancement of the escape rate due to quantum fluctuations. Eq. (17) can be expressed in terms of $\omega_{0}$ and $\omega_{b}$ as
$r_{sc}\approx r_{c}e^{\frac{\hbar\omega_{0}}{4k_{B}T}\left(1+\frac{\omega_{b}^{2}}{\omega_{0}^{2}}\right)}.$ (18)
V Discussion
The result (17) is consistent in form with those from the quantum Smoluchowski equations considered in QK1 ; A3 , except the quantum fluctuation in position $\hbar G$ is replaced by the symbol $\lambda$ in these two references. However, the viewpoints adopted in the derivations of the results are different.
In our semiclassical approach, quantum corrections are incorporated in the drift potential $V(x)$, and their values are determined by quasi-stationary conditions and minimal uncertainty relation. On the contrary, in QK1 ; A3 a quantum version of the Smoluchowski equation was deduced heuristically by modifying the diffusion coefficient $D=k_{B}T/M\gamma$ to S1 ; S2
$D_{q}=D\left[1-\frac{\lambda V^{\prime\prime}(x)}{k_{B}T}\right]^{-1},$
where $\lambda=\langle x^{2}\rangle-\langle x^{2}\rangle_{cl}$ is the quantum fluctuation in position. The values of $\lambda$ is substantiated by path-integral computations for different damping cases and temperature regimes.
In this respect we note that an earlier version of the quantum Smoluchowski equation proposed in A1 ; A2 ; S1 ; S2 , quantum corrections appeared in both the drift potential and the diffusion term. As a result, the exponent in the quantum enhancement factor in (17) is doubled. This has been noted in W2 ; QK1 ; A3 .
VI Acknowledgments
The work is supported in part by the Ministry of Science and Technology (MOST) of the Republic of China under Grants NSTC 112-2112-M-032-007 and NSTC 113-2112-M-032-010.